# Capivara: A Spectral-based Segmentation Method for IFU Data Cubes

### v1

@@ Abstract
We present capivara, a fast and scalable multi-decomposition package designed to study astrophysical properties within distinct structural components of galaxies. Our spectro-decomposition code for analyzing integral field unit (IFU) data enables a more holistic approach, moving beyond conventional radial gradients and the bulge-plus-disk dichotomy. It facilitates comprehensive comparisons of integrated stellar ages and metallicities across various galactic structures. Our classification method naturally identifies outliers and organizes the different pixels based on their dominant spectral features. The algorithm leverages the scalability and GPU acceleration of Torch, outputting both a one-dimensional spectrum and a full data cube for each galaxy component, without relying on Voronoi binning. We demonstrate the capabilities of our approach using a sample of galaxies from the MaNGA survey, processing the resulting data cubes with the starlight spectral fitting code to derive both stellar population and ionized gas properties of the galaxy components. Our method effectively groups regions with similar spectral properties in both the continuum and emission lines. By aggregating the spectra of these regions, we enhance the signal-to-noise ratio of our analysis while significantly speeding up computations by reducing the number of spectra processed simultaneously. capivara is freely available on GitHub.
@@
Data analysis-methods, galaxies: structure – galaxies: evolution![[Uncaptioned image]](extracted/5962224/capivara.jpeg)
![[Uncaptioned image]](extracted/5962224/capivara.jpeg)
@@ 1 Introduction
Galaxies, the cosmic building blocks, comprise diverse stellar components such as bulges, disks, bars, spiral arms, and nuclear star clusters, each characterized by its age, metallicity, and kinematics (e.g., Hubble, 1926; Tabor et al., 2017; Coccato et al., 2018; Johnston et al., 2022; Häußler et al., 2022; Nedkova et al., 2024; Jegatheesan et al., 2024; Lima-Dias et al., 2024; Zanatta et al., 2024).
The bulge of the Milky Way is notably ancient, exceeding 10 billion years in age, and is rich in metals. This contrasts with its disk, which exhibits metallicity closer to that of the Sun and a considerably younger average age (Hayden et al., 2015; Ness & Freeman, 2016; Chiti et al., 2021). Upon detailed analysis, the Milky Way reveals thick and thin disks, a bar, spiral arms, a pseudo-bulge, and a halo (e.g., Anders et al., 2014; Helmi, 2020; Queiroz et al., 2021; Ardern-Arentsen et al., 2024). Other models have been suggested to explain the formation of the Galactic disk, involving processes such as star formation, gas accretion, galaxy mergers, and stellar migration (Di Matteo, 2016; Imig et al., 2023).
@@
Unlike our Galaxy, such detailed scrutiny of structural components in galaxies beyond the Local Group is not possible, often leading to crude divisions between disk and bulge components (Buitrago et al., 2008; Nedkova et al., 2024). Yet, the distinction between bulge and disk populations is pivotal for understanding the processes behind galaxy formation and evolution. It is generally accepted that massive galaxies form hierarchically in a simplified two-phase process: a rapid dissipative phase (which includes a starburst phase), followed by multiple mergers (Whitney et al., 2021). The James Webb Space Telescope is showing that massive galaxies are already in place at (Carnall et al., 2023). Bulges are typically older and more metal-rich relative to disks (Lah et al., 2023). The influence of bulges and disks on galaxy evolution depends on factors like galaxy/halo mass, environment, and redshift (Conselice, 2014; Cimatti et al., 2019). A galaxy’s environment (Peng et al., 2010b) also influences its properties (Lani et al., 2013; Lima-Dias et al., 2021), with central massive galaxies typically growing inside-out due to mechanisms like gas accretion and Active Galactic Nucleus (AGN) feedback, which predominantly impact central regions (e.g., Kim et al., 2011; Krumholz et al., 2018; Baker et al., 2023). Conversely, satellite galaxies often exhibit outside-in profiles due to external quenching forces that suppress star formation from the edges inward, a phenomenon underscored by mechanisms such as major mergers (Lin et al., 2019; Cortese et al., 2021; Samuel et al., 2022). Recent findings of outside-in radial gradients in distant, low-mass satellites (Sandles et al., 2023) emphasize environmental quenching’s role from as early as . Alongside stellar histories, gas distribution and properties offer insights into quenching processes, with internal mechanisms leading to central outflows and environmental quenching leaving extended shocked gas distributions, identifiable by specific emission line ratios (Carniani et al., 2023).
Examining galaxies beyond the Local Group faces the challenge of not resolving individual stars, necessitating the use of models that approximate a galaxy’s light as stemming from a combination of single-age, single-metallicity populations (SSPs; see, e.g., Coelho et al., 2009; Coelho & Gadotti, 2011; López-Corredoira et al., 2017; Pernet et al., 2024). Analyses of these properties often use photometric decomposition techniques on broad-band galaxy images. 2D photometric decomposition codes like GIM2D (Simard et al., 2002) and GALFIT (Peng et al., 2002, 2010a) are widely used and reveal the complex process of galaxy substructure formation and the distribution of different stellar populations.
@@
The use of Integral Field Units (IFUs) has enabled more detailed studies of galaxy properties by providing simultaneous spectral and spatial information, allowing for the examination of chemical abundances and star formation histories in a spatially resolved manner. These observations have become an integral component of many low- and high-redshift surveys, including SAURON (Bacon et al., 2001), the Spectroscopic Imaging survey in the Near-infrared with SINFONI (SINS; Förster Schreiber et al., 2009), Mapping Nearby Galaxies at Apache Point Observatory (MaNGA; Bundy et al., 2015), Calar Alto Legacy Integral Field Area (CALIFA; Sánchez et al., 2012), and Middle Ages Galaxy Properties with Integral Field Spectroscopy (MAGPI; Foster et al., 2021).
@@
IFU data have been used in bulge-disk decomposition to identify spaxels belonging to bulge- or disk-dominated regions, enabling the examination of the properties of the stellar populations in these regions. For example, Wisotzki et al. (2003) used 2D modeling techniques on IFU data to deblend the quadruple QSO and gravitational lens HE 0435-1223. A similar approach was used to separate the nuclear and host galaxy of the central region of 3C 120 (Sánchez et al., 2004) and to extract the deblended spectra of the galaxies in the core of Abell 2218 (Sánchez et al., 2007). A noteworthy package is the Bulge-Disk Decomposition of IFU data (BUDDI; Johnston et al., 2017), which utilizes GalfitM (Häußler et al., 2013), a modified version of Galfit (Peng et al., 2002, 2010a), to model the light profile of multi-wavelength images of a galaxy. Méndez-Abreu et al. (2019) employed C2D, a multi-component decomposition, which consists of fitting a galaxy’s surface-brightness distribution at each wavelength (quasi-monochromatic image) via a 2D photometric decomposition code.
@@
One limitation of current image decomposition methods is their reliance on preset categories like halos, bulges, disks, bars, and rings, leaving out structures that don’t fit these categories, such as asymmetric features. Additionally, when working with IFUs, these methods typically treat each wavelength independently rather than as part of a high-dimensional tensor, potentially overlooking important correlations—such as those between bluer wavelengths and H i emission. This results in the loss of useful information that could provide insights into the physical mechanisms behind these features.
@@
This work explores an alternative approach for spectral categorization of different physical regions in galaxies. Instead of using preconceived 2D profile models, our method employs a non-parametric, multi-component decomposition based solely on spectral similarity. This allows for greater flexibility in decomposing complex structures, such as those found in mergers and irregular galaxies. By utilizing unsupervised hierarchical clustering, our approach directly targets the wavelength space instead of treating the data cube as a stack of images at each wavelength. This methodology leverages torch (Paszke et al., 2019), a deep learning library designed for tensor computations and GPU acceleration, enabling fast and efficient matrix calculations. Its capabilities make it particularly suitable for the decomposition of large datasets, offering a powerful tool for the analysis of IFU data cubes.
@@
2 Method: Spectro-Decomposition
Our model classifies each IFU pixel into clusters, ensuring that spectra within a cluster are similar, while those in different clusters are dissimilar. Here, denotes the total number of pixels , and is the number of spectral channels. We measure similarity using the distance metric, given by:
where represents the Euclidean distance. Here, and represent two independent vectors of features, is the index of the feature, and is the number of features. In our case, the features are the flux values at each wavelength bin of the IFU. While other possible choices exist, such as the Kullback–Leibler divergence (Kullback & Leibler, 1951), mutual information (Li, 1990), and so forth, we opt for the norm due to its simplicity and its relation to statistics, a standard method for spectral fitting and template matching. This approach preserves the common underlying assumption that spectra can be compared via their respective Euclidean distances. The properties of the spectra include their overall shape, the strength of particular spectral lines, and other distinguishing characteristics that can influence the similarity measure.
@@ 4 Analysis of five MaNGA galaxies
We demonstrate the feasibility of our method for identifying regions with different spectral properties within an extended object by selecting a sample of five morphologically diverse galaxies from MaNGA. The main properties of each galaxy, as listed in Table 1, were obtained from NED111http://ned.ipac.caltech.edu/ and Marvin222https://dr17.sdss.org/marvin/. Our sample includes spiral, elliptical, and merger galaxies, as well as various features such as bars, rings, and individual stars in the field.
@@
After running capivara on the MaNGA data cubes, we obtained a matrix for each galaxy, assigning each pixel to its respective group. Fig 2 shows, in the top row, a composite image of the five chosen galaxies in our sample. In the bottom row, we show the spatial distribution of the 20 components detected by capivara. The color scheme is qualitative since the components are not ordered in any continuous scale333The color palette for this figure was created using Vincent van Gogh’s The Starry Night painting as inspiration.
@@
4.1 Continuum classification
Stellar population synthesis involves several steps: first, the observed galaxy spectrum is cleaned and calibrated to remove noise and errors. The spectrum is then divided into wavelength bins, and flux is measured in each bin. Finally, the observed spectrum is fitted with theoretical templates derived from galaxy evolution models, stellar libraries, and initial mass functions to analyze the continuum.
@@ 4.1 Continuum classification
The fitting process minimizes the between observed and modeled spectra, considering factors like dust reddening and kinematics. Once the best-fit template is identified, the galaxy’s physical properties can be inferred. For instance, redshift is determined from the shift in spectral features, while luminosity is derived from the overall flux. The star formation history is deduced from the strengths of various absorption and emission lines. By evaluating the contributions of different templates, we can also infer properties such as age, and metallicity. Prior to the fitting, spurious features and emission lines are masked to ensure a more accurate study on the stellar continuum.
@@
We employed the starlight code (Cid Fernandes et al., 2004, 2005), which fits observed spectra using a model library that accounts for kinematics and dust reddening. We utilized the E-MILES library of SSP models (Vazdekis et al., 2016), based on a revised Kroupa (2001) initial mass function and BaSTI isochrones (Pietrinferni et al., 2004, 2006). This library spans from 1680 to 50,000 Å, with an optical resolution of 2.5 Å. starlight has been rigorously tested across multiple IFS galaxy surveys; for an application to MaNGA data, see Mallmann et al. (2018). Since capivara outputs are in FITS format, any spectral fitting method can easily be applied to extract stellar and ionized-gas properties from our galaxy structures.
@@
To expedite our analysis, we restricted our library to 10 representative ages4440.03, 0.06, 0.1, 0.3, 0.6, 1.0, 3.0, 6.0, 10.0, and 13.0 Gyrs and 3 representative metallicities555Z = 0.0040, 0.0198, and 0.0400 . We also included a featureless continuum (FC) with F to account for potential AGN emission, as well as very young ( 5 Myr) SPs, which are indistinguishable from AGN-type continua (Storchi-Bergmann et al., 2000; Riffel et al., 2009). Caution is therefore advised when interpreting this component. Figure 3 shows the integrated observed spectra of each galaxy in black, with the corresponding starlight model spectra in red.
@@
Since small differences in stellar population composition are often obscured by noise present in real data (Cid Fernandes et al., 2004, 2005; Riffel et al., 2009), we grouped the SSP models into four age bins: young (xy: 100 Myr), young-intermediate (xyi: 0.1- 0.7 Gyr), intermediate-old (xio: 0.7-2 Gyr), and old (xo: 2-13 Gyr). This categorization helps capture major evolutionary phases while accounting for the noise limitations in the data.
@@
In Figs. 4 to 5, we present the light fraction of the FC for each age bin. The top row shows the results before decomposition, while the bottom row displays the results after decomposition. Our method effectively identifies the main regions in each galaxy, preserving their distinct features. For the galaxy 7443-12703, our code successfully identified two regions with younger stellar populations and stronger FC contributions: one concentrated in the nucleus and another located 40 East and 80 North of the nucleus. Additionally, it accurately grouped the semi-ring structure observed in the age bin panel. In the galaxy 8135-12701, our method identified the bulge and disk regions, with the bulge dominated by an old stellar population (SP) and the disk primarily by intermediate-age SPs. Additionally, it detected a spiral arm with higher FC contributions compared to the rest of the galaxy, visible in the upper left region of 8135-12701 FC panel. This FC component likely originates from very young SPs in the spiral arms, as there is no evidence of AGN activity in this galaxy.
@@
Galaxy 8443-6102, the only early-type galaxy in our sample, exhibits the simplest stellar population structure. It consists of an older central region and outer areas within our Field of View (FoV), where a small FC contribution (10%) was detected. Additionally, a subtle ring-like structure with a slightly younger stellar population is present. Our method accurately identified all these regions, particularly the faint ring-like structure, which is barely visible in the old age bin panel. Unlike 8443-6102, 10224-6104 displays a complex structure with three distinct light centers and an uneven stellar population. The rightmost light source is dominated by an FC-type continuum, the leftmost by a young stellar population, and the central, brightest source by a young-intermediate stellar population. capivara was able to detect each light source and group their influence regions together, highlighting these differences. Additionally, a fourth region with a different continuum composition was detected in the left edge of our FoV, showing a slightly higher FC contribution if compared to its surroundings.

### v2

@@ Abstract
[…]
@@
Data analysis-methods, galaxies: structure – galaxies: evolution
Capivara: A Spectral-based Segmentation Method for IFU Data Cubes–
Capivara: A Spectral-based Segmentation Method for IFU Data Cubes
@@ 1 Introduction
Galaxies, the fundamental building blocks of the cosmos, comprise a rich diversity of stellar components, including bulges, disks, bars, spiral arms, and nuclear star clusters. Each of these structures exhibits distinct signatures in age, metallicity, and kinematics, shaped by the interplay of galactic formation and evolution over cosmic time (e.g., Hubble, 1926; Cappellari, 2016; Tabor et al., 2017; Coccato et al., 2018; Sánchez, 2020; Johnston et al., 2022; Häußler et al., 2022; Nedkova et al., 2024; Jegatheesan et al., 2024; Lima-Dias et al., 2024; Zanatta et al., 2024). The bulge of the Milky Way primarily consists of an ancient population over 10 billion years old and rich in metals, but it also contains intermediate-age, thick-disk-like stars and a relatively young nuclear disk. This contrasts with its thin disk, which exhibits metallicity closer to that of the Sun and a considerably younger average age (Hayden et al., 2015; Howes et al., 2015; Ness & Freeman, 2016; Chiti et al., 2021).
@@
Upon detailed analysis, the Milky Way reveals thick and thin disks, a bar, spiral arms, a nuclear disk, and a complex halo (e.g., Anders et al., 2014; Helmi, 2020; Queiroz et al., 2021; Ardern-Arentsen et al., 2024). Other models have been suggested to explain the formation of the Galactic disk, involving processes such as star formation, gas accretion, galaxy mergers, and stellar migration (Di Matteo, 2016; Imig et al., 2023). Unlike our Galaxy, such detailed scrutiny of structural components in galaxies beyond the Local Group is not possible, often leading to crude divisions between disk and bulge components (Buitrago et al., 2008; Nedkova et al., 2024), although advances have been achieved for nearby galaxies with newer instrumentation (Gadotti et al., 2019; van de Sande et al., 2024). Yet, the distinction between bulge and disk populations is pivotal for understanding the processes behind galaxy formation and evolution. It is generally accepted that massive galaxies form hierarchically in a simplified two-phase process: a rapid dissipative phase (which includes a starburst phase), followed by multiple mergers (Whitney et al., 2021). The James Webb Space Telescope is showing that massive galaxies are already in place at (Carnall et al., 2023). Bulges are typically older and more metal-rich relative to disks (Lah et al., 2023). The influence of bulges and disks on galaxy evolution depends on factors like galaxy/halo mass, environment, and redshift (Conselice, 2014; Cimatti et al., 2019).
A galaxy’s environment (Peng et al., 2010b) also influences its properties (Lani et al., 2013; Lima-Dias et al., 2021), with central massive galaxies typically growing inside-out due to mechanisms like gas accretion and Active Galactic Nucleus (AGN) feedback, which predominantly impact central regions (e.g., Croton et al., 2006; Goddard et al., 2017; Kim et al., 2011; Krumholz et al., 2018; Baker et al., 2025). Conversely, satellite galaxies often exhibit outside-in profiles due to external quenching forces that suppress star formation from the edges inward, a phenomenon underscored by mechanisms such as major mergers (Lin et al., 2019; Cortese et al., 2021; Samuel et al., 2022). Recent findings of outside-in radial gradients in distant, low-mass satellites (Sandles et al., 2023) emphasize environmental quenching’s role from as early as . Alongside stellar histories, gas distribution and properties offer insights into quenching processes, with internal mechanisms leading to central outflows and environmental quenching leaving extended shocked gas distributions, identifiable by specific emission line ratios (Carniani et al., 2024).
@@
The use of Integral Field Units (IFUs) has enabled more detailed studies of galaxy properties by providing simultaneous spectral and spatial information, allowing for the examination of chemical abundances, kinematics, and star formation histories in a spatially resolved manner. These observations are an integral part of numerous low- and high-redshift surveys, including SAURON (Bacon et al., 2001), the Spectroscopic Imaging Survey in the Near-infrared with SINFONI (SINS; Förster Schreiber et al., 2009), Mapping Nearby Galaxies at Apache Point Observatory (MaNGA; Bundy et al., 2015), Calar Alto Legacy Integral Field Area (CALIFA; Sánchez et al., 2012), the Sydney Australian Astronomical Observatory Multi-Object Integral Field Spectrograph (SAMI; Croom et al., 2012, 2021) Galaxy Survey, the Gemini Near-Infrared Field Spectrograph (NIFS; Riffel et al., 2018), Time Inference with MUSE in Extragalactic Rings (TIMER; Gadotti et al., 2019), and the Middle Ages Galaxy Properties with Integral Field Spectroscopy (MAGPI; Foster et al., 2021).
@@
IFU data have been used in morphological decomposition to identify spaxels belonging to bulge- or disk-dominated regions (and other structures), enabling the examination of the properties of the stellar populations in these regions. For example, Wisotzki et al. (2003) used 2D modeling techniques on IFU data to deblend the quadruple QSO and gravitational lens HE 0435-1223. A similar approach was used to separate the nuclear and host galaxy of the central region of 3C 120 (Sánchez et al., 2004) and to extract the deblended spectra of the galaxies in the core of Abell 2218 (Sánchez et al., 2007). A noteworthy package is the Bulge-Disk Decomposition of IFU data (BUDDI; Johnston et al., 2017), which utilizes GalfitM (Häußler et al., 2013), a modified version of Galfit (Peng et al., 2002, 2010a), to model the light profile of multi-wavelength images of a galaxy. Méndez-Abreu et al. (2019) employed C2D, a multi-component decomposition, which consists of fitting a galaxy’s surface-brightness distribution at each wavelength (quasi-monochromatic image) via a 2D photometric decomposition code. Bittner et al. (2021) studies the kinematic and stellar population of inner bars in three disk galaxies, based on the multi-component photometric decompositions from de Lorenzo-Cáceres et al. (2019).
@@
One shortcoming of current image decomposition methods is their reliance on preset categories—such as halos, bulges, disks, bars, and rings—which exclude structures that do not fit, like asymmetric features. Additionally, when working with IFUs, these methods typically treat each wavelength independently rather than as part of a high-dimensional tensor, potentially overlooking important correlations—such as those between bluer wavelengths and H ii emission. This results in the loss of useful information that could provide insights into the physical mechanisms behind these features.
@@
This work explores an alternative approach for spectral categorization of different physical regions in galaxies. Instead of using preconceived two-dimensional profile models, our method employs a spectral-based segmentation approach solely relying on spectral similarity. This allows for greater flexibility in characterizing complex structures, such as those found in mergers and irregular galaxies. By utilizing unsupervised hierarchical clustering, our approach directly targets the wavelength space instead of treating the data cube as a stack of images at each wavelength. This methodology leverages torch (Paszke et al., 2019), a deep learning library designed for tensor computations and GPU acceleration, enabling fast and efficient matrix calculations. Its capabilities make it particularly suitable for the segmentation of large datasets, offering a powerful tool for the analysis of IFU data cubes.
@@
2 Method: Spectro-Segmentation
Our model classifies each IFU pixel into clusters, ensuring that spectra within a cluster are similar, while those in different clusters are dissimilar. Here, denotes the total number of pixels , and is the number of spectral wavelengths. We measure similarity using the distance metric, given by:
where represents the Euclidean distance. Here, and represent two independent vectors of features, is the index of the feature, and is the number of features.
In our case, the features are the flux values at each wavelength bin of a spaxel in the IFU. While other possible choices exist, such as the Kullback–Leibler divergence (Kullback & Leibler, 1951), mutual information (Li, 1990), and so forth, we opt for the norm due to its simplicity and its relation to statistics, a standard method for spectral fitting and template matching. This approach preserves the common underlying assumption that spectra can be compared via their respective Euclidean distances. The properties of the spectra include their overall shape, the strength of particular spectral lines, and other distinguishing characteristics that can influence the similarity measure.
@@ 4 Analysis of five MaNGA galaxies
We demonstrate the feasibility of our method for identifying regions with different spectral properties within an extended object by selecting a sample of five morphologically diverse galaxies from MaNGA. The main properties of each galaxy, as listed in Table 1, were obtained from NED111http://ned.ipac.caltech.edu/ and Marvin222https://dr17.sdss.org/marvin/ (Aguado et al., 2019). Our sample includes spiral, elliptical, and merger galaxies, as well as various features such as bars, rings, and individual stars in the field.
@@
After applying the capivara algorithm to the MaNGA data cubes, we obtained segmentation matrices assigning each pixel of the galaxy images to their respective component groups. Fig. 2 presents SDSS composite images of five selected galaxies from our sample in the top row, with their MaNGA plate numbers indicated in the upper-right corner and the corresponding IFU fields of view outlined in purple. The middle row displays the spatial distributions of the 20 components detected by capivara, while the bottom row shows the spatial distributions obtained using the vorbin package for the same galaxies. Comparing the segmentation maps from both approaches provides insight into their sensitivity in capturing distinct morphological and structural features. The color schemes employed are qualitative, as the components represent discrete groups without an implied continuous ordering.333The color palette employed was inspired by Vincent van Gogh’s painting The Starry Night (González-Gaitán et al., 2019).
@@
In Fig. 3 we compare our results with Voronoi tessellation at various granularity levels. We run vorbin across a range of target signal-to-noise (S/N) ratios: 50, 100, 150, 200, 250. Each target S/N corresponds inversely to the number of resulting regions—the higher the target S/N, the fewer the regions generated. Subsequently, we run capivara with a comparable number of regions to facilitate a meaningful comparison between the two segmentation methods. capivara preserves an aesthetically pleasing visual resemblance to the composite SDSS image. This visual coherence is desirable, as our objective extends beyond mere S/N enhancement—we aim to group pixels into physically meaningful regions reflective of intrinsic galactic structures.
Finally, to provide both a qualitative and quantitative impression of our model compared to standard Voronoi binning, we conducted an illustrative experiment with six groups. This choice, made without loss of generality, serves merely to convey the intuitive qualitative differences between the two strategies. A key desideratum for any clustering approach is to enhance the SNR while preserving spectral similarity. Increasing the SNR by combining highly dissimilar spectra may indeed improve noise characteristics; however, this also dilutes the intrinsic physical properties of the region444For example, reaching the target SNR may require summing a large number of spaxels, but this predominantly accumulates noise, thereby merging regions that are not physically connected.. Figure 4 illustrates this point. In both the top and bottom maps, each color-coded region corresponds to a cluster of pixels found by capivara and vorbin, respectively. On the right, we show all spectra associated with a given region, annotated with two key metrics: n, the number of pixels in the cluster, and the average pairwise correlation among all spectra within that cluster (AvgCorr). Higher values of AvgCorr indicate more homogeneous spectral shapes, reflecting the internal consistency of each cluster. By comparing the left and right panels, one can observe both the spatial distribution of each cluster and the degree to which those pixels share a common spectral signature—highlighting capivara’s ability to preserve shape information throughout the data.
@@
4.1 Stellar continuum classification
Examining galaxies beyond the Local Group presents the challenge of not resolving individual stars, which necessitates modeling a galaxy’s integrated light (e.g., colors or full spectrum) as a combination of simple stellar populations (SSPs; see, e.g., Coelho & Gadotti, 2011; López-Corredoira et al., 2017; Bittner et al., 2020; Pernet et al., 2024). These combinations are compared with the galaxy’s observed emission, and inversion techniques are employed to infer its physical properties. In many cases, this process begins with photometric decomposition applied to broad-band galaxy images.
@@ 4.1 Stellar continuum classification
Spectral fitting, on the other hand, involves matching observed spectra with SSP models while accounting for physical effects such as dust attenuation and Doppler shifts. A common approach is the inversion technique, in which a linear combination of SSP models is fitted to the observed spectrum, incorporating processes like reddening and stellar velocity dispersion. The best fit is typically determined by minimizing the statistic (Walcher et al., 2011). Once the best-fit template is identified, the galaxy’s physical properties—such as age, metallicity, and star formation history—can be inferred. The star formation history can be deduced from the light contribution of stellar population models in each age bin to the best-fit template. By averaging the contributions of the different SSPs, we can also infer integrated properties such as light- and mass-weighted ages and metallicities. Prior to the fitting, spurious features and emission lines are often masked to ensure a more accurate study on the stellar continuum.
@@
We employed the starlight code (Cid Fernandes et al., 2004, 2005), which fits observed spectra using a model library that accounts for kinematics and dust reddening. We utilized the E-MILES library of SSP models (Vazdekis et al., 2016), based on a revised Kroupa (2001) initial mass function and BaSTI isochrones (Pietrinferni et al., 2004, 2006). This library spans from 1680 to 50,000 Å, with an optical resolution of FWHM = 2.5 Å. starlight has been rigorously tested across multiple IFS galaxy surveys; for an application to MaNGA data, see Mallmann et al. (2018). Since capivara outputs are in FITS format, any spectral fitting method can easily be applied to extract stellar and ionized-gas properties from our galaxy structures.
@@
To expedite our analysis, we restricted our library to 10 representative ages5550.03, 0.06, 0.1, 0.3, 0.6, 1.0, 3.0, 6.0, 10.0, and 13.0 Gyrs and 3 representative metallicities666Z = 0.2, 1, and 2 . We also included a featureless continuum (FC) with F to account for potential AGN emission, as well as very young ( 5 Myr) SSPs, which are indistinguishable from AGN-type continua (Koski, 1978; Storchi-Bergmann et al., 2000; Riffel et al., 2009). Caution is therefore advised when interpreting this component. Figure 5 shows the integrated observed spectra of each galaxy in black, with the corresponding starlight model spectra in red777For galaxy 10224-6104, the adopted SSP base may not fully capture its stellar populations. Incorporating younger populations reveals clear signs of recent star formation. See https://manga.if.ufrgs.br/explorer/215 for an interactive visualization.
@@
Since small differences in stellar population composition are often obscured by noise in real data (Cid Fernandes et al., 2004, 2005; Riffel et al., 2009), we grouped the SSP models into three age bins: young (xy: 100 Myr), intermediate-age (xi: 0.1–2.0 Gyr), and old (xo: 2–13 Gyr). This categorization reflects the dominant contributors to the integrated light in each bin: blue, hot stars in the young population; evolved giants, particularly TP-AGB stars, in the intermediate-age range; and cooler, red stars in the old population.
@@
In Figs. 6 to 7, we present FC and condensed stellar population vectors in light fraction. The top row shows the results before decomposition, while the bottom row displays the results after decomposition. Our method effectively identifies the main regions in each galaxy, preserving their distinct features. For the galaxy 7443-12703, our code successfully identified two regions with younger stellar populations and stronger FC contributions: one concentrated in the nucleus and another located 40 East and 80 North of the nucleus. Additionally, it accurately grouped the semi-ring structure observed in the age bin panel. In the galaxy 8135-12701, our method identified the bulge and disk regions, with the bulge dominated by an old stellar population (SP) and the disk primarily by intermediate-age SPs. Additionally, our method identified a spiral arm with elevated FC contributions compared to the rest of the galaxy, clearly visible in the upper left region of the 8135-12701 FC panel. This apparent FC excess is likely a consequence of fitting degeneracies, where starlight is driven to compensate for limitations in the base set’s representation of the youngest populations. The fitting process is influenced by reddening and flux calibration effects, which often force unphysical negative values that artificially shift the fits toward bluer solutions (see, e.g., Riffel et al., 2022, for a more detailed discussion).
@@
Galaxy 8443-6102, the only early-type galaxy in our sample, exhibits the simplest stellar population structure. It consists of an older central region and outer areas within our Field of View (FoV), where a small FC contribution (10%) was detected. Additionally, a subtle ring-like structure with a slightly younger stellar population is present. Our method accurately identified all these regions, particularly the faint ring-like structure, which is barely visible in the old age bin panel. Unlike 8443-6102, 10224-6104 displays a complex structure with three distinct light centers and an uneven stellar population. The rightmost light source is dominated by an FC-type continuum, the leftmost by a young stellar population, and the central, brightest source by an intermediate-age stellar population. capivara was able to detect each light source and group their influence regions together, highlighting these differences. Additionally, a fourth region with a different continuum composition was detected in the left edge of our FoV, showing a slightly higher FC contribution if compared to its surroundings.